# Asymptotics for conformal inference

### v1

@@ Abstract
Conformal inference is a versatile tool for building prediction sets in regression or classification. In this paper, we consider the false coverage proportion (FCP) in a transductive setting with a calibration sample of $n$ points and a test sample of $m$ points. We identify the exact, distribution-free, asymptotic distribution of the FCP when both $n$ and $m$ tend to infinity. This shows in particular that FCP control can be achieved by using the well-known Kolmogorov distribution, and puts forward that the asymptotic variance is decreasing in the ratio $n/m$. We then provide a number of extensions by considering the novelty detection problem, weighted conformal inference and distribution shift between the calibration sample and the test sample. In particular, our asymptotical results allow to accurately quantify the asymptotical behavior of the errors when weighted conformal inference is used.
@@ 1.1 Background
In classical statistics, producing prediction sets for outcomes often relies on strong model assumptions. Recent advances involve complex data sets and sophisticated machine learning methods, for which such an approach is not appropriate. One recent solution is conformal prediction (Saunders et al.,, 1999; Vovk et al.,, 2005; Angelopoulos and Bates,, 2021) which consists in calibrating the prediction set according to an appropriate quantile of a calibration/training sample. Strikingly, this method provides a finite-sample valid coverage (that is, for any size $n\geqslant 1$ of the calibration sample), for any underlying distribution of the data and for any underlying point-prediction machine learning algorithm. Similar techniques can be employed for the novelty detection task (Balasubramanian et al.,, 2014; Bates et al.,, 2023; Marandon et al.,, 2024).
@@ 1.2 Aim and contributions
We consider here the so-called transductive setting (Vovk,, 2013), where it is given a calibration sample of $n$ points $(X_{1},Y_{1})$, $\dots$, $(X_{n},Y_{n})$ and a test sample of $m$ points $(X_{n+1},Y_{n+1}),$dots, $(X_{n+m},Y_{n+m})$. While the calibration sample is fully observed, the $Y_{i}$’s of the test points are not observed and a prediction set should be provided for each of them. The false coverage proportion (FCP) for the $m$ conformal prediction sets $\mathcal{C}^{\alpha}(X_{n+1}),\dots,\mathcal{C}^{\alpha}(X_{n+m})$ (see below for a formal definition) is given as the proportion of coverage errors among the test sample:
@@
Under standard assumptions, the distribution of the process $\mathrm{FCP}_{m}^{(n)}$ has been shown to be distribution-free, in the sense that it does only depend on $n$ and $m$ (Marques F.,, 2023; Huang et al.,, 2024; Gazin et al.,, 2024). Due to the dependence between the individual coverage errors, this distribution is particularly complex in $n$ and $m$ and combinatorial formulas have been derived in Marques F., (2023); Huang et al., (2024); Gazin et al., (2024). Nevertheless, focusing on the maximum absolute deviation
@@
with $I_{n}(\alpha)=\lfloor(n+1)\alpha\rfloor/(n+1)$, a DKW-type concentration inequality has been derived in Gazin et al., (2024), which is both simple and finite-sample valid. It explicitly involves a rate $\tau_{n,m}^{1/2}$ defined by
$\tau_{n,m}\coloneqq\frac{nm}{n+m}\in\left[\frac{n\wedge m}{2},n\wedge m\right].$ (2)
However, this DKW inequality is conservative in general (see Figure 1 below), which makes the corresponding FCP control conservative.
The aim of this paper is to complement the above studies by analyzing $\mathrm{FCP}_{m}^{(n)}$ from an asymptotical point of view, where both $m$ and $n$ tend to infinity. Our contributions are as follows:
@@
    We show that $\mathrm{FCP}_{m}^{(n)}$ converges uniformly to the nominal value at rate $\tau_{n,m}^{1/2}$ and that the asymptotic covariance process is a standard Brownian bridge (Theorem 3.1). Compared to the “oracle” case where $n=\infty$, it means that the variance is inflated by a factor asymptotically equivalent to $(n+m)/n$, for instance $2$ in the case where $n\sim m$.
@@
    A direct corollary of this result is that $\tau_{n,m}^{1/2}\|\mathrm{FCP}_{m}^{(n)}-I_{n}\|_{\infty}$ converges to the well known Kolmogorov distribution, that is, the distribution with c.d.f. $x\mapsto(1-2\sum_{k\geqslant 1}(-1)^{k}e^{-2k^{2}x^{2}}){\mathds{1}_{x\geqslant 0}}$. A comparison between the quantiles of this distribution and those given by empirical simulations or DKW is provided in Figure 1. As we can see, while the validity of the new quantile is only asymptotic with $n,m\to+\infty$, the asymptotic quantile is simple and more accurate than the quantile obtained from DKW. Hence, the new result allows to get simple, accurate and asymptotically-valid confidence bounds for the FCP process and more generally one can have an asymptotic approximation of all quantities related to the distribution of the $\mathrm{FCP}$ process.
@@
    We then extend this result to the case where the distribution of the calibration sample is not equal to the test sample, that is, under a distribution shift (Theorem 3.2). As expected, the convergence of the FCP is not towards the nominal level in this case but rather towards a new term $G$ that takes into account this shift. The asymptotic covariance process is also modified according to $G$ and an explicit formula is given.
@@
    To recover the appropriate nominal level in the limit, we adopt the weighted conformal approach (Tibshirani et al.,, 2019; Barber et al.,, 2023), with specific weights that rely on the data distribution, that we refer to as oracle weights. The central limit theorem shares similarities with the exchangeable case described above, with the essential difference that the asymptotic covariance process is not distribution free, and depends on the sample distributions (Theorem 3.4).
@@
    We also obtained a convergence result in case of non-oracle weights (Theorem 3.3), which is crucial to quantify the FCP asymptotic behavior in the difficult but realistic case where the user has not access to the true distribution shift. An illustration in displayed in Figure 2, where the asymptotical confidence interval of the FCP is given in function of an error $\Delta$, measuring how the used weights deviate from the oracle ones. This puts forward that the FCP gets significantly away from $\alpha$ when $\Delta$ is above $\approx 0.133$ or below $\approx-0.11$ in this framework.
@@
    Finally, we obtain similar results for the novelty detection task, by studying the asymptotical behavior of the false discovery proportion (FDP) of classical procedures (Bates et al.,, 2023; Jin and Candès,, 2023), see Section 4. To our knowledge, using a weighted approach has not been considered before in the novelty detection case.
The proofs are based on specific decompositions of the processes that can be found in Section 5, while further details are postponed to appendices.
Figure 1: Comparison of the $(1-\delta)$-quantile of different approximations of the distribution of $\|\mathrm{FCP}_{m}^{(n)}-I_{n}\|_{\infty}$ in (1) for different values of $n,m$ and $\delta$. The approximations include Monte-Carlo ($1000$ replications), DKW (Gazin et al.,, 2024) and the new asymptotic one (see text).
Figure 2: Plot of the asymptotical confidence interval for the $\mathrm{FCP}$ of the weighted conformal method (at level $80\%$) obtained in Theorem 3.3 versus an error parameter $\Delta$. The calibration sample and the test sample are distributed according to the exponential distribution with mean $1$ and $1/3$, respectively. The weight function used in the conformal method is $w_{\Delta}(x)=\exp(-(2+\Delta)x){\mathds{1}_{x>0}}$, which corresponds to the oracle choice if $\Delta=0$ (for which the asymptotic average of $\mathrm{FCP}(\alpha)$ is equal to $\alpha$) and deviates from it if $\Delta\neq 0$.
@@ 1.3 Relation to previous work
Conformal prediction is a general pipeline and we refer the reader to Vovk et al., (2005) or Angelopoulos and Bates, (2021) for reviews. We focus here on the inductive/split conformal inference (Papadopoulos et al.,, 2002), where an independent training sample is used to build the predictors, while a calibration sample is used to adjust the prediction sets. All our results can be thought of holding conditionally on the training sample, that is, they hold once the point-predictions have been computed.
@@
The process $\mathrm{FCP}_{m}^{(n)}$ coincides with the empirical cumulative distribution function (e.c.d.f.) of conformal $p$-values as introduced by Saunders et al., (1999). As recalled above, for a given nominal level $\alpha$, the non-asymptotic distribution of $\mathrm{FCP}_{m}^{(n)}(\alpha)$ has been obtained in Marques F., (2023); Huang et al., (2024) and the full distribution of the process $\mathrm{FCP}_{m}^{(n)}$ has been given in Gazin et al., (2024). In addition, Marques F., (2023) gives an asymptotic result for $\mathrm{FCP}_{m}^{(n)}(\alpha)$ in the regime where $n$ tends to infinity only after having made $m$ tends to infinity, whereas here we make $n$ and $m$ grow to infinity independently. In addition, our results are uniform in $\alpha$. Next, in Nguyen et al., (2024), they study a more general risk and also obtained asymptotical results, but only as $m$ tends to infinity while $n$ is kept fixed.
@@
Finally, we study here novelty detection procedures obtained by applying the Benjamini-Hochberg procedure ($\mathrm{BH}$, Benjamini and Hochberg,, 1995) to the conformal $p$-values (Mary and Roquain,, 2022; Bates et al.,, 2023; Marandon et al.,, 2024) or to the weighted conformal $p$-values (Tibshirani et al.,, 2019; Barber et al.,, 2023; Jin and Candès,, 2023). Asymptotics for the FDP of such procedures have been extensively studied in the literature, see Genovese and Wasserman, (2002); Neuvial, (2008) for iid uniform $p$-values and Delattre and Roquain, (2011, 2016); Kluger and Owen, (2024) for several types of dependence structures. The present work is in this line of research by seeing (weighted) conformal $p$-values as a particular case of dependent $p$-values, with a very specific dependence structure, induced by the calibration sample. We obtain our results in the novelty detection setting by using the techniques introduced in Neuvial, (2008), that combine functional central limit theorems with the functional delta method. In particular, this provides the full asymptotic FDP distribution for the procedure of Jin and Candès, (2023) (or a variant thereof), for which only in-expectation results were known to our knowledge.
@@ 2.1 Prediction setting
We consider the classical split/inductive conformal prediction (Papadopoulos et al.,, 2002). A calibration set $\{Z_{k},k\in\llbracket n\rrbracket\}$ is observed with $Z_{k}=(X_{k},Y_{k})$ and, given a new point $Z_{n+1}=(X_{n+1},Y_{n+1})$ for which only the covariate $X_{n+1}\in\mathcal{X}$ is observed, a prediction set should be inferred for the outcome $Y_{n+1}\in\mathcal{Y}$ [^1] The regression and classification settings correspond to $\mathcal{Y}=\mathbb{R}$ and $\mathcal{Y}$ finite, respectively.. A non-conformity score function $S:\mathcal{X}\times\mathcal{Y}\rightarrow\mathbb{R}$ is also given; $S(X,Y)\in\mathbb{R}$ measures the non-conformity of the response $Y\in\mathcal{Y}$ with the covariate $X\in\mathcal{X}$. The classical example in regression is the residual $S(X,Y)=|Y-{\widehat{\mu}}(X)|$ where ${\widehat{\mu}}:\mathcal{X}\rightarrow\mathcal{Y}$ is a regression function trained from an independent training sample (considered as fixed here). The (split) conformal prediction set at level $(1-\alpha)$ for $X_{n+1}$, denoted by $\mathcal{C}^{\alpha}(X_{n+1})$, is defined as
@@
$\mathcal{C}^{\alpha}(X_{n+1})\coloneqq\left\{y\in\mathcal{Y}:\ S(X_{n+1},y)\leqslant S_{\left(\lceil(n+1)(1-\alpha)\rceil\right)}\right\},$ (3)
where $S_{(1)}\leqslant S_{(2)}\leqslant\cdots\leqslant S_{(n)}<S_{(n+1)}:=+\infty$ correspond to the ordered calibration scores $\mathcal{D}_{{\tiny\mbox{cal}}}=\left\{S_{k}:{k\in\llbracket n\rrbracket}\right\}$. The set $\mathcal{C}^{\alpha}(X_{n+1})$ can be equivalently described as follows (this classical fact can be retrieved from Lemma F.1): $Y_{n+1}\notin\mathcal{C}^{\alpha}(X_{n+1})$ if and only if ${p_{1}^{(n)}}\leqslant\alpha$, where the conformal $p$-value is given by
@@
${p_{1}^{(n)}}\coloneqq(n+1)^{-1}\left(1+\sum_{k\in\llbracket n\rrbracket}{\mathds{1}_{S_{k}\geqslant T_{1}}}\right),$
where $T_{1}=S(X_{n+1},Y_{n+1})$.
In this paper, we consider a transductive setting (Vovk,, 2013), where a decision should be made for a whole test sample of size $m$ for which only the covariates are observed. We denote $\mathcal{D}_{{\tiny\mbox{test}}}=(T_{i},\ i\in\llbracket m\rrbracket)$ the set of the unobserved (since we only observed the covariates of the test points) scores of the test sample. Considering the conformal prediction sets $\mathcal{C}^{\alpha}(X_{n+i})$, $i\in\llbracket m\rrbracket$ — corresponding to (3) for all members of the test sample — gives rise to a family of $m$ conformal $p$-values $({p_{i}^{(n)}})_{i\in\llbracket m\rrbracket}$ defined by
@@
${p_{i}^{(n)}}\coloneqq(n+1)^{-1}\left(1+\sum_{k\in\llbracket n\rrbracket}{\mathds{1}_{S_{k}\geqslant T_{i}}}\right),\>\>\>i\in\llbracket m\rrbracket.$ (4)
@@
The false coverage proportion $\mathrm{FCP}_{m}^{(n)}(\alpha)$ of the prediction set family $\left(\mathcal{C}^{\alpha}(X_{n+i})\right)_{i\in\llbracket m\rrbracket}$ is defined by
@@
Note that this corresponds to the empirical cumulative density function (e.c.d.f.) of the conformal $p$-values family $({p_{i}^{(n)}})_{i\in\llbracket m\rrbracket}$. The following assumption will be considered throughout the paper for the (weighted or not) conformal prediction task.
@@
The set of calibration scores $\left(S_{k}\right)_{k\geqslant 1}$ and the set of test scores $\left(T_{i}\right)_{i\geqslant 1}$ are two independent families of real random variables. The $S_{k},k\geqslant 1,$ (resp. $T_{i},i\geqslant 1$) are i.i.d. with distribution $P_{\tiny\mbox{cal}}$ (resp. $P_{\tiny\mbox{test}}$) and cumulative distribution function $F_{\tiny\mbox{cal}}$ (resp. $F_{\tiny\mbox{test}}$). Moreover, $F_{\tiny\mbox{cal}}$ and $F_{\tiny\mbox{test}}$ are continuous functions.

### v3

@@ Abstract
Conformal inference is a versatile tool for building prediction sets in regression or classification. We study the false coverage proportion (FCP) in a simultaneous inference setting with a calibration sample of $n$ points and a test sample of $m$ points. We identify the exact, distribution-free, asymptotic distribution of the FCP when both $n$ and $m$ tend to infinity. This shows in particular that FCP control can be achieved by using the well-known Kolmogorov distribution, and puts forward that the asymptotic variance is decreasing in the ratio $n/m$. We then provide a number of extensions by considering the problems of novelty detection, weighted conformal inference or distribution shift between the calibration sample and the test sample. In particular, our asymptotic results allow to accurately quantify the asymptotic behavior of the errors (a miscovering interval or declaring a false novelty) when weighted conformal inference is used.
weighted conformal prediction,
keywords:
@@ 1.1 Background
In classical statistics, producing prediction sets for outcomes often relies on strong model assumptions. Recent advances involve complex data sets and sophisticated machine learning methods, for which such an approach is not appropriate. One recent solution is given by conformal prediction (Saunders, Gammerman and Vovk, 1999; Vovk, Gammerman and Shafer, 2005; Angelopoulos and Bates, 2021) which consists in calibrating the prediction set according to an appropriate quantile of a calibration/training sample. Strikingly, this method provides a finite-sample valid coverage (i.e. valid for any size $n\geqslant 1$ of the calibration sample) for any underlying distribution of the data and for any underlying point-prediction machine learning algorithm. Similar techniques can be employed for the novelty detection task (Balasubramanian, Ho and Vovk, 2014; Bates et al., 2023; Marandon et al., 2024).
@@ 1.2 Aim and contributions
We consider here the multiple setting (Vovk, 2013), where it is given a training dataset (which is considered as fixed here), a calibration sample of $n$ points $(X_{1},Y_{1})$, $\dots$, $(X_{n},Y_{n})$ and a test sample of $m$ points $(X_{n+1},Y_{n+1}),$ $\dots$, $(X_{n+m},Y_{n+m})$. While the calibration sample is fully observed, the $Y_{i}$’s of the test points are not observed and a prediction set should be provided for each of them. The false coverage proportion (FCP) for the $m$ conformal prediction sets $\mathcal{C}^{\alpha}(X_{n+1}),\dots,\mathcal{C}^{\alpha}(X_{n+m})$ (see below for a formal definition) is defined as the proportion of coverage errors among the test sample:
@@
Under standard assumptions and when the conformal prediction sets come from the inductive/split conformal procedure the distribution of the process $\mathrm{FCP}_{m}^{(n)}$ has been shown to be distribution-free, in the sense that it does only depend on $n$ and $m$ (Marques F., 2025; Huang et al., 2024; Gazin, Blanchard and Roquain, 2024). Due to the dependence between the individual coverage errors, combinatorial formulas derived in those references are particulary complex in $m$ and $n$. Nevertheless, focusing on the maximum absolute deviation
@@
with $I_{n}(\alpha)=\lfloor(n+1)\alpha\rfloor/(n+1)$, a DKW-type concentration inequality has been derived in Gazin, Blanchard and Roquain (2024), which is both simple and finite-sample valid. It explicitly involves a rate $\tau_{n,m}^{1/2}$ with
$\tau_{n,m}\coloneqq\frac{nm}{n+m}\in\mathopen{}\mathclose{{\left[\frac{n\wedge m}{2},n\wedge m}}\right].$ (2)
However, this DKW inequality is conservative in general (see Figure 1 below), which makes the corresponding FCP control stringent.
The aim of this paper is to complement the above studies by analyzing $\mathrm{FCP}_{m}^{(n)}$ from an asymptotic point of view, where both $m$ and $n$ tend to infinity which is equivalent to the convergence of $\tau_{n,m}$ to infinity. This asymptotic study will also cover the case of a distribution shift i.e. when the exchangeability assumption does not hold thus extending the analysis of Gazin, Blanchard and Roquain (2024) and more broadlythe usual framework of conformal inference. We also study, in the same asymptotic regime, the asymptotic of the Benjamini-Hochberg procedure when conformal $p$-values are used (Bates et al., 2023), which complement the non-asymptotic work of Bates et al. (2023); Gazin, Blanchard and Roquain (2024) among others. We refer to Section 1.3 for more details on the relation of this work with previous studies. More precisely, our contributions are the following ones:
@@
    We show that under standard assumptions $\mathrm{FCP}_{m}^{(n)}$ converges uniformly to the nominal value at rate $\tau_{n,m}^{1/2}$ and that the asymptotic covariance process is a standard Brownian bridge (Theorem 3.1). Compared to the “oracle” case where $n=\infty$, it means that the variance is inflated by a factor asymptotically equivalent to $(n+m)/n$, for instance $2$ in the case where $n\sim m$.
@@
    A direct corollary of this result is that $\tau_{n,m}^{1/2}\|\mathrm{FCP}_{m}^{(n)}-I_{n}\|_{\infty}$ converges to the well known Kolmogorov distribution, that is, the distribution with c.d.f. $x\mapsto(1-2\sum_{k\geqslant 1}(-1)^{k}e^{-2k^{2}x^{2}}){\mathds{1}_{x\geqslant 0}}$. A comparison between the quantiles of this distribution, those given by empirical simulations of $\mathrm{FCP}_{m}^{(n)}$ or those from the Conformal-DKW is provided in Figure 1. While its validity is only asymptotic with $n,m\to+\infty$, it appears that the Kolmogorov quantile is simple and more accurate than the quantile obtained from DKW. Hence, our limit theorem allows to get simple, accurate and asymptotically-valid confidence bounds for the FCP process as well as asymptotic approximations of all quantities related to the distribution of the $\mathrm{FCP}$ process.
@@
    We then extend this result to the case where the distribution of the calibration sample is not equal to the test sample, that is, under a distribution shift (Theorem 3.2). As expected, the convergence of the FCP is not towards the nominal level in this case but rather towards a new term $G$ that takes this shift into account. The asymptotic covariance process is also modified according to $G$ and an explicit formula is given. The role of the ratio $n/(n+m)$ (number of calibration points divided by the total number of points) clearly appears in the limit and indicates the existence of three asymptotic regimes: under, proportional and over calibrated, see Table 1.
@@
    To recover the appropriate nominal level in the limit, we adopt the weighted conformal approach (Tibshirani et al., 2019; Barber et al., 2023) with specific weights that rely on the data distribution, that we refer to as oracle weights. The obtaned central limit theorem shares similarities with the exchangeable case described above, with the essential difference that the asymptotic covariance process is not distribution free, and depends on the sample distributions (Theorem 3.4).
@@
    We also obtain a convergence result in case of non-oracle weights (Theorem 3.3), which is crucial to quantify the FCP asymptotic behavior in the difficult but realistic case where the user has not access to the true distribution shift. An illustration is displayed in Figure 2, where the asymptotic confidence interval for the FCP is given as a function of an error $\Delta$, measuring how the used weights deviate from the oracle ones. This puts forward that the FCP gets significantly away from $\alpha$ when $\Delta$ is above $\approx 0.133$ or below $\approx-0.11$ in this framework.
@@
    Finally, we obtain similar results for the novelty detection task, by studying the asymptotic behavior of the false discovery proportion (FDP) of classical procedures (Bates et al., 2023; Jin and Candès, 2023) (Theorems 4.1 and 4.2).
The proofs are based on specific decompositions of the processes that can be found in Section 5, while further details are postponed to appendices, notably Section G with a presentation of the main results of the paper under relaxed assumptions.
Figure 1: Comparison of the $(1-\delta)$-quantile of different approximations of the distribution of $\|\mathrm{FCP}_{m}^{(n)}-I_{n}\|_{\infty}$ in (1) for different values of $n,m$ and $\delta$. The approximations include Monte-Carlo (10001000 replications), Conformal-DKW (Gazin, Blanchard and Roquain, 2024) and the asymptotic exact Kolmogorov one.
Figure 2: Plot of the asymptotic confidence interval for the $\mathrm{FCP}$ of the weighted conformal method (at level $80\%$) obtained in Theorem 3.3 versus a parameter estimation error $\Delta$. The calibration sample and the test sample are distributed according to the exponential distribution with mean $1$ and $1/3$, respectively. The weight function used in the conformal method is $w_{\Delta}(x)=\exp(-(2+\Delta)x){\mathds{1}_{x>0}}$, which corresponds to the oracle choice if $\Delta=0$ (for which the asymptotic average of $\mathrm{FCP}(\alpha)$ is equal to $\alpha$) and deviates from it if $\Delta\neq 0$.
@@ 1.3 Relation to previous work
Conformal prediction is a general pipeline and we refer the reader to Vovk, Gammerman and Shafer (2005) or Angelopoulos and Bates (2021) for reviews. We focus here on the inductive/split conformal inference (Papadopoulos et al., 2002), where an independent training sample is used to build the predictors, while a calibration sample is used to adjust the prediction sets. All our results can be thought of holding conditionally on the training sample, that is, they hold once the point-predictions have been computed.
@@
The process $\mathrm{FCP}_{m}^{(n)}$ coincides with the empirical cumulative distribution function (e.c.d.f.) of conformal $p$-values as introduced by Saunders, Gammerman and Vovk (1999). As recalled above, for a given nominal level $\alpha$, the non-asymptotic distribution of $\mathrm{FCP}_{m}^{(n)}(\alpha)$ has been obtained in Marques F. (2025); Huang et al. (2024) and the full distribution of the process $\mathrm{FCP}_{m}^{(n)}$ has been given in Gazin, Blanchard and Roquain (2024). The main difference with the current study is that they provide non-asymptotic properties under the exchangeable setting while we provide asymptotic properties under the exchangeable and the non-exchangeable frameworks. In addition, Marques F. (2025) gives an asymptotic result for $\mathrm{FCP}_{m}^{(n)}(\alpha)$ with $\alpha\in(0,1)$ fixed in the regime where $n$ tends to infinity only after making $m$ tend to infinity, whereas here we make $n$ and $m$ grow to infinity independently leading to a double asymptotic. Furthermore all our results are uniform in $\alpha$.
Next, the asymptotic distribution of $\mathrm{FCP}_{m}^{(n)}$ with $n$ fixed and $m\rightarrow+\infty$ is also studied with a PAC point of view since this limit is equal to the miscoverage probability given the calibration sample $\mathbb{P}[Y_{n+1}\notin\mathcal{C}^{\alpha}(X_{n}+1)|(X_{i},Y_{i})_{i\in\llbracket n\rrbracket}]$: Vovk (2012) derives Hoeffding-like tail bounds for a fixed $\alpha$, and other bounds was derived by Bian and Barber (2022); Sarkar and Kuchibhotla (2023) the latter being uniform in $\alpha$. Angelopoulos and Bates (2021) states that $\mathbb{P}[Y_{n+1}\notin\mathcal{C}^{\alpha}(X_{n}+1)|(X_{i},Y_{i})_{i\in\llbracket n\rrbracket}]$ follows a Beta distribution, which is exactly the limit identified by Marques F. (2025). Similarly, Nguyen et al. (2024) obtained asymptotic results for the error as $m$ tends to infinity while $n$ is kept fixed but with more general risk function and not necessary the 010-1 loss function. In the novelty detection setting, Bates et al. (2023) also provided bounds for the type I error given the calibration sample uniformly in $\alpha$.
@@
Finally, we study here novelty detection procedures obtained by applying the Benjamini-Hochberg procedure ($\mathrm{BH}$, Benjamini and Hochberg, 1995) to the conformal $p$-values (Mary and Roquain, 2022; Bates et al., 2023; Marandon et al., 2024) or to the weighted conformal $p$-values (Tibshirani et al., 2019; Barber et al., 2023; Jin and Candès, 2023). Asymptotics for the FDP of such procedures have been extensively studied in the literature, see Genovese and Wasserman (2002, 2004); Neuvial (2008) for i.i.d. uniform $p$-values and Delattre and Roquain (2011, 2016); Kluger and Owen (2024) for several types of dependence structures. The present work is in this line of research by seeing (weighted) conformal $p$-values as a particular case of dependent $p$-values having a specific dependency structure induced by the common calibration sample. In particular, our work shows that the ratio $n/(n+m)$ measures the dependency strength in the double asymptotic regime where both $n$ (calibration sample size) and $m$ (the number of test) tend to infinity. In the novelty detection setting, we derive our results by using the techniques introduced in Neuvial (2008), that combine functional central limit theorems with the functional delta method. In particular, we provide the full asymptotic FDP distribution for the procedure of Jin and Candès (2023) (or a variant thereof), for which only in-expectation results were known.
@@ 2.1 Prediction setting
We consider the classical inductive/split conformal prediction (Papadopoulos et al., 2002). A calibration set $\{Z_{k},k\in\llbracket n\rrbracket\}$ is observed with $Z_{k}=(X_{k},Y_{k})$ and, given a new point $Z_{n+1}=(X_{n+1},Y_{n+1})$ for which only the covariate $X_{n+1}\in\mathcal{X}$ is observed, a prediction set should be inferred for the outcome $Y_{n+1}\in\mathcal{Y}$. [^1] The regression and classification settings correspond to $\mathcal{Y}=\mathbb{R}$ and $\mathcal{Y}$ finite, respectively. A training dataset $\mathcal{D}_{{\tiny\mbox{train}}}$ independent of the test point and the calibration sample is observed and used to compute a non-conformity score function $S:\mathcal{X}\times\mathcal{Y}\rightarrow\mathbb{R}$. The scalar $S(X,Y)\in\mathbb{R}$ measures the non-conformity of the response $Y\in\mathcal{Y}$ with the covariate $X\in\mathcal{X}$. All our results hold conditionally on the training dataset, hence to alleviate notation we will consider the score function $S$ as fixed and determinist. This is in accordance with the usual split conformal model: the training/calibration split is considered as already done, with a fixed independent training sample and a growing calibration sample. A classical example in regression is the absolute value of the residual $S(X,Y)=|Y-{\widehat{\mu}}(X)|$ where ${\widehat{\mu}}:\mathcal{X}\rightarrow\mathcal{Y}$ is a regression function trained from an independent training sample (considered as fixed here). The (split) conformal prediction set at level $(1-\alpha)$ for $X_{n+1}$, denoted by $\mathcal{C}^{\alpha}(X_{n+1})$, is defined as
@@
$\mathcal{C}^{\alpha}(X_{n+1})\coloneqq\mathopen{}\mathclose{{\left\{y\in\mathcal{Y}:\ S(X_{n+1},y)\leqslant S_{\mathopen{}\mathclose{{\left(\lceil(n+1)(1-\alpha)\rceil}}\right)}}}\right\},$ (3)
where $S_{(1)}\leqslant S_{(2)}\leqslant\cdots\leqslant S_{(n)}<S_{(n+1)}:=+\infty$ correspond to the ordered calibration scores $\mathcal{D}_{{\tiny\mbox{cal}}}\coloneqq\mathopen{}\mathclose{{\left\{S_{k}:{k\in\llbracket n\rrbracket}}}\right\}=\mathopen{}\mathclose{{\left\{S(X_{k},Y_{k})\,:\,k\in\llbracket n\rrbracket}}\right\}$. The interest of the conformal prediction set is the validity of the prediction region under weak assumption. More precisely, if the random variables $(Z_{1},\cdots,Z_{n},Z_{n+1})$ (i.e. the $n$ calibration points and the new test point) are exchangeable (or i.i.d.) then we have the marginal coverage property: $\mathbb{P}(Y_{n+1}\notin\mathcal{C}^{\alpha}(X_{n+1}))\leqslant\alpha$. Furthermore, if the scores $(S(X_{i},Y_{i}))_{i\in\llbracket n+1\rrbracket}$ are exchangeable and without ties almost surely, then $\mathbb{P}(Y_{n+1}\notin\mathcal{C}^{\alpha}(X_{n+1}))$ is lower bounded by $\alpha-(n+1)^{-1}$ (see Vovk, Gammerman and Shafer, 2005 for more details). The set $\mathcal{C}^{\alpha}(X_{n+1})$ can be equivalently described by using $p$-values (this classical fact can be retrieved from Lemma F.1). In particular, by defining the (unobserved) conformal $p$-value $p_{1}^{(n)}$} as in (4) with the test score $T_{1}=S(X_{n+1},Y_{n+1})$ (see below), we obtain that $Y_{n+1}\notin\mathcal{C}^{\alpha}(X_{n+1})$ if and only if ${p_{1}^{(n)}}\leqslant\alpha$. By the above marginal coverage property, $p_{1}^{(n)}$} is a valid $p$-value for testing exchangeability of $(Z_{1},\cdots,Z_{n},Z_{n+1})$ since it is super-uniformly distributed under exchangeability. In this paper, we consider a multiple/simultaneous inference setting (also called transductive by Vovk, 2013; Gazin, Blanchard and Roquain, 2024; Behboodi et al., 2025), where a decision should be made for a whole test sample of size $m$ $(Z_{n+1},\dots,Z_{n+m})$ with $Z_{n+i}=(X_{n+i},Y_{n+i})$, for which only the covariates $X_{n+1},\dots,X_{n+m}$ are observed. We denote $\mathcal{D}_{{\tiny\mbox{test}}}\coloneqq\mathopen{}\mathclose{{\left\{T_{i}\,:\,i\in\llbracket m\rrbracket}}\right\}$ with $T_{i}=S(X_{n+i},Y_{n+i})$ the set of the unobserved (since only the $X_{n+i}$’s are observed) scores of the test sample. Considering the conformal prediction sets $\mathcal{C}^{\alpha}(X_{n+i})$, $i\in\llbracket m\rrbracket$ — corresponding to (3) for any test point — gives rise to a family of $m$ conformal $p$-values $({p_{i}^{(n)}})_{i\in\llbracket m\rrbracket}$ defined by
@@
${p_{i}^{(n)}}\coloneqq(n+1)^{-1}\mathopen{}\mathclose{{\left(1+\sum_{k\in\llbracket n\rrbracket}{\mathds{1}_{S_{k}\geqslant T_{i}}}}}\right),\>\>\>i\in\llbracket m\rrbracket.$ (4)
We also define the family of theoretical $p$-values $({p_{i}^{(+\infty)}})_{i\in\llbracket m\rrbracket}$ when the calibration scores are i.i.d. as
@@
${p_{i}^{(+\infty)}}\coloneqq\mathbb{P}\mathopen{}\mathclose{{\left[S_{1}\geqslant T_{i}\middle|T_{i}}}\right],\>\>\>i\in\llbracket m\rrbracket.$ (5)
@@
By applying the law of large number when the calibration scores are i.i.d., the $p$-value $p_{i}^{(+\infty)}$} can be seen as the limit of $p_{i}^{(n)}$} when $n$ tends to infinity. The false coverage proportion $\mathrm{FCP}_{m}^{(n)}(\alpha)$ of the prediction set family $\mathopen{}\mathclose{{\left(\mathcal{C}^{\alpha}(X_{n+i})}}\right)_{i\in\llbracket m\rrbracket}$ is defined by
@@
Note that this corresponds to the empirical cumulative distribution function (e.c.d.f.) of the conformal $p$-values family $({p_{i}^{(n)}})_{i\in\llbracket m\rrbracket}$. We emphasise that the set $\mathcal{D}_{{\tiny\mbox{test}}}$ of test scores is not observed, and neither are the conformal $p$-values (4) in the prediction setting. The $\mathrm{FCP}$ is not observed, and the aim is to study its behavior when the sizes of the calibration and test samples both tend to infinity. The following assumption will be considered throughout the paper for the (weighted or not) conformal prediction task.
@@
The set of calibration scores $\mathopen{}\mathclose{{\left(S_{k}}}\right)_{k\geqslant 1}$ and the set of test scores $\mathopen{}\mathclose{{\left(T_{i}}}\right)_{i\geqslant 1}$ are two independent families of real random variables. The $S_{k}$’s ($k\geqslant 1$), (resp. $T_{i}$’s ($i\geqslant 1$)) are i.i.d. with distribution $P_{\tiny\mbox{cal}}$ (resp. $P_{\tiny\mbox{test}}$) and cumulative distribution function $F_{\tiny\mbox{cal}}$ (resp. $F_{\tiny\mbox{test}}$). Moreover, $F_{\tiny\mbox{cal}}$ and $F_{\tiny\mbox{test}}$ are continuous functions.